### v3

# On the Analysis of Misspecified Variational Inequalities with Nonlinear Constraints

Abstract
In this paper, we study a class of misspecified variational inequalities (VIs) where both the monotone operator and nonlinear convex constraints depend on an unknown parameter learned via a secondary VI. Existing data-driven VI methods typically follow a decoupled learn-then-optimize scheme, causing the approximation error from the learning to propagate the main decision-making problem and hinder convergence. We instead consider a simultaneous approach that jointly solves the main and secondary VIs. To efficiently handle nonlinear constraints with parameter misspecification, we propose a single-loop inexact Augmented Lagrangian method that simultaneously updates the primal decision variables, dual multipliers, and the misspecified parameter. The method combines a forward-reflected-backward step with an Augmented Lagrangian penalty, and explicitly handles misspecification on both the operator and constraint functions. Moreover, we introduce a relaxed performance metric based on the Minty VI gap combined with an aggregated infeasibility metric. By proving boundedness of the dual iterates, we establish $\mathcal{O}(1/K)$ ergodic convergence rates for these metrics. Numerical Experiments are provided to showcase the superior performance of our algorithm compared to state-of-the-art methods.
Keywords: Variational inequalities, Augmented Lagrangian method, Misspecified and data-driven optimization
1 Introduction
Variational inequality (VI) problems have received significant attention in recent years owing to their general formulation that unifies a wide spectrum of well-known problems, including constrained and unconstrained optimization, saddle point (SP) problems, and Nash equilibrium problems [8]. These formulations naturally arise in diverse domains such as machine learning, economics, and game theory. Classical formulations assume complete knowledge of the problem parameters and feasible sets, enabling deterministic analysis of equilibria and their stability properties. However, in many practical applications, ranging from networked systems and energy markets to learning-based decision models, the underlying parameters governing the VI are not known precisely and must instead be inferred from data or auxiliary learning processes. This leads to misspecified variational inequalities, where the true problem data depend on an unknown parameter vector that is estimated imperfectly. Such misspecification introduces an inherent coupling between the estimation and optimization components, giving rise to new analytical and computational challenges.
In contrast to stochastic or robust optimization approaches [11, 15, 12, 13] that assume access to distributional information or uncertainty sets, misspecified VIs capture settings in which the unknown parameters evolve through a separate, possibly data-driven, mechanism. This perspective bridges classical equilibrium modeling with modern learning paradigms, allowing for adaptive decision-making under uncertainty without explicit probabilistic assumptions. While recent advances have established foundational results for misspecified VIs Ahmadi and Shanbhag [3], Jiang and Shanbhag [10] in unconstrained or affine settings, the incorporation of nonlinear constraints remains largely unexplored.
In this paper, we address a class of misspecified constrained variational inequality in which both the operator and the constraint functions depend on an unknown parameter vector $\theta^{*}$. Specifically, we seek $(x^{*},\theta^{*})$ such that
$F(x^{*},\theta^{*})^{\top}(y-x^{*})\geq 0,\quad\forall y\in\mathcal{X}(\theta^{*}),$ (Misspecified VI)
where $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is a continuous mapping, and $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the feasible set defined by
$\mathcal{X}(\theta^{*})\triangleq\left\{x\in X\mid f_{j}(x,\theta^{*})\leq 0,\quad\forall j=1,\dotsc,J\right\},\quad\textit{ and }\quad X\triangleq\prod_{i=1}^{N}X_{i}.$ (1)
Each $X_{i}\subseteq\mathbb{R}^{n_{i}}$ is a nonempty, closed, and convex set, and $f_{j}(\cdot,\theta^{*})$ is convex in $x$. The parameter $\theta^{*}\in\Theta\subseteq\mathbb{R}^{m}$ is not known a priori; instead, it is determined as the solution of a secondary variational inequality
$H(\theta^{*})^{\top}(\vartheta-\theta^{*})\geq 0,\quad\forall\vartheta\in\Theta,$ (2)
where $H:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is a continuous and strongly monotone mapping, and $\Theta\subseteq\mathbb{R}^{m}$ is a nonempty, closed, and convex set. This coupled VI-within-VI framework simulates, for instance, network or market equilibria where the feasible set and operator are contingent upon demand or cost parameters derived from historical data, or policy-constrained Cournot competition with unobserved prices collected from empirical observations. Our focus is on algorithmic frameworks that simultaneously update the primal variable $x$, the dual multipliers associated with the nonlinear constraints, and the parameter estimate $\theta$. To address the challenges posed by nonlinear constraints in the misspecified VI framework, we employ an augmented Lagrangian approach [5]. This method embeds the nonlinear constraints into the objective through a Lagrange multiplier term combined with a quadratic penalty, thereby enforcing feasibility progressively.
1.1 Literature Review
A wide body of work has examined VI problems, including foundational results for deterministic VIs [20, 16, 17, 18] and their stochastic counterparts [12, 15, 14]. In this section, we review prevailing methodologies for optimization problems involving unknown or misspecified parameters, studied across frameworks such as constrained minimization, saddle point formulations, and VI models.
Constraint-based Minimization. In this line of research, numerous works focus on objective function minimization under both deterministic [1, 2, 6] and stochastic frameworks [10, 23, 21]. In the deterministic setting, [2] study a misspecified convex optimization problem of the form $\min_{x\in X}f(x;\theta^{*})$ subject to $\theta^{*}\in\arg\min_{\theta\in\Theta}g(\theta)$. When both the optimization and learning objectives are strongly convex, the coupled update scheme achieves a linear convergence rate. In contrast, when $f(\cdot;\theta^{*})$ is merely convex while $g$ remains strongly convex and a weak–sharpness condition is satisfied, the standard $\mathcal{O}(1/K)$ rate is retained. [6] consider a convex optimization problem with a constraint set that depends on an unknown parameter and introduce an inexact parametric augmented Lagrangian method (IPALM) to handle this misspecification. The scheme employs an accelerated proximal-gradient step for updating the optimization decision variable and an outer update of the multipliers. They derive iteration-complexity bounds showing that, under constant and increasing penalty sequences, IPALM requires at most $\mathcal{O}(K^{-1/4})$ and $\mathcal{O}(\log K/K)$, respectively.
Saddle Point (SP) and Variational Inequality (VI) Problems. Recent research has extended the study of misspecification beyond classical optimization to SP and VI problems. In [3], a class of misspecified monotone VI problems is studied, where misspecifications may arise in the objective. To tackle these issues, the authors develop extragradient and regularized first-order algorithms tailored for misspecified monotone VI problems coupled with strongly convex learning subproblems, thereby guaranteeing stable convergence. Complementary to this, [10] analyze stochastic variational inequality (SVI) problems in the context of misspecification, covering both stochastic convex optimization and a range of stochastic equilibrium models. Their analysis critically relies on the strong convexity of the learning problem, which guarantees convergence of both the optimization and learning iterates. They further establish almost-sure convergence in strongly monotone settings, achieving the optimal rate of $\mathcal{O}(1/K)$, and in merely monotone settings by employing iterative Tikhonov regularization, which yields a rate of $\mathcal{O}(\sqrt{\log K}/\sqrt{K})$ under weak sharpness. [3] extend their earlier work [2] to misspecified monotone variational inequality problems and show that a constant–steplength misspecified extragradient scheme achieves asymptotic convergence of the iterates. More recently, Ahmadi and Hamedani [4] examine a class of misspecified convex-concave SP problems where the misspecified parameter can be obtained through a secondary SP problem with a strongly convex-concave objective function. They introduce a Learning-aware Accelerated Primal-Dual method that integrates the parameter updates directly into the dual momentum step and utilizes an adaptive backtracking scheme to select step sizes adaptively. Their approach attains a convergence rate of $\mathcal{O}(\log K/K)$ without relying on compactness of the feasible sets. In addition, they extend their method to settings where the secondary SP problem may not have a unique solution, proving that a modified version of the algorithm tailored to this structure achieves a rate of $\mathcal{O}(1/\sqrt{K})$ in terms of both gap function and learning suboptimality.
1.2 Research Gap and Our Contribution
Previous work on misspecified VIs has primarily considered settings where the constraint set is easy to project onto and does not depend on any unknown parameters. In contrast, this paper focuses on a broader class of misspecified VIs that involve misspecified nonlinear constraints in the primary problem. A detailed comparison between our method and existing approaches is provided in Table 1.
More specifically, we study a coupled VI framework in which the primary VI (Misspecified VI) is defined over a parameter-dependent feasible region $\mathcal{X}(\theta)$, specified by a set of convex nonlinear functions. The unknown parameter $\theta^{*}$ is simultaneously estimated by solving a secondary VI problem (2), which we assume to be strongly monotone. To address this problem, we propose a novel single-loop method that combines augmented Lagrangian and forward-reflected-backward step. To contend with the challenge of misspecified constraints, we show that under suitable assumptions, the corresponding dual iterates remain bounded, which helps us to establish the convergence rate of the proposed method. Furthermore, to measure the quality of the solution, we introduce a relaxed Minty VI gap as well as infeasibility metrics and establish an ergodic convergence rate of $\mathcal{O}(1/K)$ in a monotone setting, matching the optimal first-order convergence rate for fully specified monotone VIs [12, 15].
Table 1: Comparison of related misspecified convex optimization/monotone VI studies.
[…]
Ahmadi and Shanbhag [3] VI VI -
Aybat et al. [6] Optimization Optimization $\mathcal{O}(\frac{\log K}{K})$
Jiang and Shanbhag [10] Stochastic VI Stochastic VI $\mathcal{O}\left(\frac{\sqrt{\log K}}{\sqrt{K}}\right)$
Ahmadi and Hamedani [4] Saddle Point Saddle Point $\mathcal{O}(\frac{\log K}{K})$
This work VI VI $\mathbf{\mathcal{O}(\frac{1}{K})}$
1.3 Motivating Example
The Misspecified constrained VIs in problem (Misspecified VI) have broad applications in markets and networked systems where feasibility depends on misspecified parameters learned from data. Some of these examples include constrained Nash-Cournot competition characterized by demand misspecification and policy cap [9], inverse and data-driven equilibrium in [19], misspecified Markowitz portfolio optimization problem in [6], and minimax schemes with expectation constraints [22]. Next, we provide a Cournot example in the following motivating example.
Cournot with a price-function constraint and unknown slope:
Modern commodity markets often operate under policy rules that constrain prices while firms face uncertainty about demand responsiveness. Consider a homogeneous goods market with $N$ firms, $i=1,\dots,N$. Firm $i$ chooses a quantity $x_{i}\geq 0$; let the aggregate output be $X\triangleq\sum_{i=1}^{N}x_{i}$. The market clears at a linear inverse demand $p(X;b^{*})=a^{*}-b^{*}X,$ where $a^{*}>0$ is known and $b^{*}>0$ is the unknown slope capturing how aggressively price falls with output. Firms know costs and capacities but do not know the true $b^{*}$. Beyond firm-level capacity limits $x_{i}\in[0,\mathrm{Cap}_{i}]$, many markets also face a price-function constraint induced by policy, such as an administrative price cap. We encode this as $p(X;b^{*})\leq a^{*}-\delta$, for some $\delta>0$. Mathematically, it imposes a lower bound on the feasible output that depends on the (unknown) slope $b^{*}$. Let $c_{i}:[0,\mathrm{Cap}_{i}]\to\mathbb{R}_{+}$ be firm $i$’s convex cost. Then, each firm’s objective is to maximize the profit $\pi_{i}({\bf x})\triangleq p(X;b^{*})x_{i}-c_{i}(x_{i})$ such that $x_{i}\in[0,\mathrm{Cap}_{i}]$ and $p(X;b^{*})\leq a^{*}-\delta$. In practice, the firms observe sample pairs $\{(p_{\text{obs},t},X_{t})\}_{t=1}^{T}$ and must learn $b^{*}$ from data while computing equilibria that respect the price constraint.
Formulation as a Misspecified Variational Inequality:
We cast the problem as finding a pair $(x^{*},b^{*})$ in the form of (Misspecified VI) where the secondary VI (2) represents learning parameter $b^{*}$ from $\{(p_{\text{obs},t},X_{t})\}_{t=1}^{T}$. For instance, one can use least-squares regression to learn the unknown parameter $b^{*}$ as follows:
$b^{*}=\operatorname*{\arg\!\min}_{b\geq 0}\ \mathcal{L}(b)=\frac{1}{2}\sum_{t=1}^{T}\big(p_{\text{obs},t}-(a^{*}-bX_{t})\big)^{2}.$ (3)
Moreover, the feasible set in (Misspecified VI) can be characterized as $\mathcal{X}(b^{*})\triangleq\Bigl\{\mathbf{x}=[x_{i}]_{i=1}^{N}\in\prod_{i=1}^{N}[0,\text{Cap}_{i}]\Big|p(X;b^{*})\leq a^{*}-\delta\Bigr\}$. Then, the operator $F(\mathbf{x},b)$ is defined by concatenating the gradients of the negative profit functions of the firms $F_{i}(\mathbf{x},b)\triangleq c_{i}^{\prime}(x_{i})+b(X+x_{i})-a^{*}$.
Outline of the paper:
The remainder of the paper is organized as follows. In Section 2, we introduce notations and state the assumptions. Section 3 presents the proposed algorithm along with augmented-Lagrangian formulation of the misspecified variational inequality. In Section 4, we establish convergence theory for the proposed method and derive the explicit convergence rate. Section 5 implements a misspecified Cournot equilibrium and empirically compares the proposed scheme against two benchmark methods. Detailed proofs of the lemmas and the theorem are provided in the Appendix.
2 Preliminaries and Assumptions
Notation.
Throughout the paper, $\|\cdot\|$ denotes the Euclidean norm, and for any scalar $x\in\mathbb{R}$ we use $x]_{+}\triangleq\max\{x,0\}$. For a differentiable vector-valued function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$, we write $\mathbf{J}f:\mathbb{R}^{n}\to\mathbb{R}^{m\times n}$ for its Jacobian matrix. For a closed convex set $X$ the projection operator is denoted by $\Pi_{X}(x)\triangleq\operatorname*{\arg\!\min}_{y\in X}\|x-y\|$.
We make the following standard assumptions regarding Misspecified VI problem.
Assumption 2.1.
The following hold:
  1. (i)
    For any $\theta\in\Theta$, the mapping $F(\cdot,\theta):\mathbb{R}^{n}\to\mathbb{R}^{n}$ is continuous and monotone on $X$, i.e.,
    $\left(F(x,\theta)-F(y,\theta)\right)^{\top}(x-y)\geq 0,\quad\forall x,y\in X.$
  2. (ii)
    The mapping $H:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is Lipschitz continuous and strongly monotone on $\Theta$, i.e.,
    $\|H(\theta)-H(\vartheta)\|\leq L_{H}\|\theta-\vartheta\|,\quad\left(H(\theta)-H(\vartheta)\right)^{\top}(\theta-\vartheta)\geq\mu_{H}\|\theta-\vartheta\|,\quad\forall\theta,\vartheta\in\Theta.$
  3. (iii)
    For any $j\in\{1,\ldots,J\}$ and $\theta\in\Theta$, the function $f_{j}(\cdot,\theta)$ is convex and continuously differentiable on $X$. Moreover, there exists $L_{\nabla f,x}\geq 0$ such that for all $x,y\in X$ and all $\theta\in\Theta$,
    $\|\nabla_{x}f_{j}(x,\theta)-\nabla_{x}f_{j}(y,\theta)\|\leq L_{\nabla f,x}\|x-y\|.$
    In addition, for any $\theta\in\Theta$ and $x\in X$, $f_{j}(\cdot,\theta)$ and $f_{j}(x,\cdot)$ are Lipschitz continuous with constants $L_{f,x}\geq 0$ and $L_{f,\theta}\geq 0$ respectively.
  4. (iv)
    The mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is Lipschitz continuous in $x$ and $\theta$ over $X\times\Theta$; that is, there exist constants $L_{F,x}\geq 0$ and $L_{F,\theta}\geq 0$ such that:
    $\|F(x_{1},\theta)-F(x_{2},\theta)\|\leq L_{F,x}\|x_{1}-x_{2}\|,\quad\forall x_{1},x_{2}\in X,\quad\forall\theta\in\Theta,$
    $\|F(x,\theta_{1})-F(x,\theta_{2})\|\leq L_{F,\theta}\|\theta_{1}-\theta_{2}\|,\quad\forall x\in X,\quad\forall\theta_{1},\theta_{2}\in\Theta.$
  5. (v)
    The sets $X\subseteq\mathbb{R}^{n}$ and $\Theta\subseteq\mathbb{R}^{m}$ are nonempty, closed, convex and compact.
  6. (vi)
    (Slater’s Condition) There exists $\hat{x}\in X$ such that: $f_{j}(\hat{x},\theta^{*})<0$, for all $j\in\{1,\dotsc,J\}$.
Definition 2.1.
(Augmented-Lagrangian Function for Misspecified VI) For $x,y\in\mathbb{R}^{n}$, $\lambda\in\mathbb{R}^{J}$, $\theta\in\mathbb{R}^{m}$, and $\rho>0$, the Augmented-Lagrangian function for the Misspecified VI problem is defined as
$\mathcal{L}_{\rho}(x,y,\theta,\lambda)\triangleq F(y,\theta)^{\top}(x-y)+\Phi_{\rho}(x,\lambda,\theta),$
where $\Phi_{\rho}(x,\lambda,\theta)\triangleq\sum_{j=1}^{J}\phi_{\rho}(f_{j}(x,\theta),\lambda_{j})$ and $\phi_{\rho}(u,v)\triangleq\begin{cases}uv+\frac{\rho}{2}u^{2},&\text{if }\rho u+v\geq 0,\\ -\frac{v^{2}}{2\rho},&\text{otherwise.}\end{cases}$
Based on the above definition, one can observe that
$\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})=\sum_{j=1}^{J}\left[\rho_{k}f_{j}(x_{k},\theta^{*})+\lambda_{k}^{(j)}\right]_{+}\nabla f_{j}(x_{k},\theta^{*}).$ (4)
Proposition 2.1.
(KKT Conditions) Suppose that a solution pair $(x^{*},\theta^{*})$ of (Misspecified VI) exists and Assumption 2.1 holds. Let $f(x,\theta^{*})\triangleq[f_{1}(x,\theta^{*}),\dots,f_{J}(x,\theta^{*})]^{\top}$, and the Jacobian matrix $\nabla f(x,\theta^{*})\triangleq[\nabla f_{1}(x,\theta^{*}),\dots,\nabla f_{J}(x,\theta^{*})]^{\top}\in\mathbb{R}^{n\times J}$. There exists a pair $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ and $\lambda^{*}\in\mathbb{R}^{J}$ satisfying the following KKT conditions:
  • (i)
    Stationarity: $0\in F(x^{*},\theta^{*})+\nabla f(x^{*},\theta^{*})^{\top}\lambda^{*}+\mathcal{N}_{X}(x^{*})\quad\textit{ and }\quad 0\in H(\theta^{*})+\mathcal{N}_{\Theta}(\theta^{*}),$ where $\mathcal{N}_{X}(x^{*})$, and $\mathcal{N}_{\Theta}(\theta^{*})$ are the normal cones of sets $X$ and $\Theta$ at $x^{*}$ and $\theta^{*}$, respectively.
  • (ii)
    Complementary Slackness: $0\leq\lambda^{*}\perp-f(x^{*},\theta^{*})\geq 0.$
  • (iii)
    Feasibility: $x^{*}\in\mathcal{X}(\theta^{*})\triangleq\{x\in X\mid f_{j}(x,\theta^{*})\leq 0,\quad\forall j=1,\dotsc,J\},\quad\textit{and }\quad\theta^{*}\in\Theta.$
Proof.
Proof. Note that any solution pair $(x^{*},\theta^{*})$ of (Misspecified VI) is a solution of the following optimization problem
$\min_{y\in X}y^{T}F(x^{*},\theta^{*})\quad\text{ s.t. }\quad f(y,\theta^{*})\leq\mathbf{0}.$ (5)
Since the Slater condition holds, the first-order KKT condition of (5) implies that there exists $\lambda^{*}\in\mathbb{R}^{J}$ satisfying conditions (i)-(iii). Moreover, the optimality condition of the secondary VI problem (2) immediately implies that $0\in H(\theta^{*})+\mathcal{N}_{\Theta}(\theta^{*})$. ∎
Lemma 2.2.
Consider (Misspecified VI) problem under Assumption 2.1 and let $(x^{*},\theta^{*},\lambda^{*})$ satisfy the KKT conditions of Proposition 2.1. Then for all $x\in\mathcal{X(\theta^{*})}$, $(x-x^{*})^{\top}F(x^{*},\theta^{*})+\sum_{j=1}^{J}\lambda^{*(j)}f_{j}(x,\theta^{*})\geq 0$.
Proof.
Proof The result follows directly from Lemma 1 in [5], which establishes the corresponding inequality under Assumption 2.1. ∎
3 Proposed Method
In this section, we present an augmented Lagrangian method for solving (Misspecified VI). Due to the presence of a nonlinear misspecified constraint, we introduce a multiplier $\lambda$ to relax this constraint, following Definition 2.1. Moreover, to handle the parameter misspecification, we simultaneously solve a secondary VI problem that generates a sequence of iterates converging to the unique optimal parameter $\theta^{*}$. The approximation of $\theta^{*}$ obtained at each iteration is then used to update both the main decision variable $x$ and the multiplier $\lambda$. To evaluate the quality of the obtained solution, we introduce a relaxed gap function. First, recall that the standard VI gap function for (Misspecified VI) is defined as $\text{Gap}(x,\theta^{*})\triangleq\sup_{y\in\mathcal{X}(\theta^{*})}{F(y,\theta^{*})^{\top}(x-y)},$ and an $\epsilon$-approximate solution satisfies $\text{Gap}(x,\theta^{*})\leq\epsilon$. This gap function is a valid optimality metric, since $\text{Gap}(x,\theta^{*})\geq 0$ for any feasible $x\in\mathcal{X}(\theta^{*})$, and $\text{Gap}(x,\theta^{*})=0$ implies that $x$ is a solution to (Misspecified VI). However, because our proposed method employs a dual multiplier to enforce constraints and accounts for the misspecified parameter $\theta$, the generated sequence may be infeasible. Consequently, the algorithm’s output $\bar{x}$ may not belong to the feasible set $\mathcal{X}(\theta^{*})$, and we might have $\text{Gap}(\bar{x},\theta^{*})<0$. To address this issue, we measure both (i) the level of infeasibility, quantified as $\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon$, and (ii) a relaxed gap function defined as $\widetilde{\text{Gap}}(x,\theta^{*})\triangleq\sup_{y\in\mathcal{X}_{\epsilon}(\theta^{*})}{F(y,\theta^{*})^{\top}(x-y)}$ where $\mathcal{X}_{\epsilon}(\theta^{*})\triangleq\{x\in X\mid\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon\}$ denotes the $\epsilon$-enlargement of the feasible set. It is straightforward to verify that $\widetilde{\text{Gap}}(x,\theta^{*})\geq 0$ for any $x\in\mathcal{X}_{\epsilon}(\theta^{*})$, and $\widetilde{\text{Gap}}(x,\theta^{*})\geq\text{Gap}(x,\theta^{*})$ for any $x$. Therefore, we define an $\epsilon$-approximate solution $\bar{x}\in X$ if
$\text{(Relaxed gap): }\widetilde{\text{Gap}}(\bar{x},\theta^{*})\leq\epsilon,\qquad\text{(Infasibility): }\mathbf{1}^{\top}[f(\bar{x},\theta^{*})]_{+}\leq\epsilon.$
Furthermore, as $\epsilon\to 0$, the point $\bar{x}$ becomes feasible, i.e., $\bar{x}\in\mathcal{X}(\theta^{*})$, which implies that $\widetilde{\text{Gap}}(x,\theta^{*})=\text{Gap}(x,\theta^{*})=0$, hence, $\bar{x}$ is a solution to (Misspecified VI).
To find such an approximate solution, we propose a novel single-loop algorithm by combining the forward-reflected-backward method with an augmented Lagrangian update. This combination enables us to effectively handle nonlinear constraints under misspecified evaluations of both the operator $F$ and the nonlinear constraint functions $f_{j}$’s. Recall the augmented Lagrangian function from Definition 2.1, then for fixed $(\theta,\lambda)$, the term $\Phi_{\rho}(x,\lambda,\theta)$ imposes penalization for violating the constraints $f_{j}(x,\theta)\leq 0$. In particular, at each iteration $k\geq 0$, the decision variable $x$ is updated using a projected forward operation applied to the augmented Lagrangian function combined with a reflected (correction) step $r_{k}=F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$ as $x_{k+1}=\Pi_{X}[x_{k}-\gamma_{k}(r_{k}+\nabla_{x}\mathcal{L}_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k}))]$. Then, the dual multiplier $\lambda_{k}$ is updated via a projected gradient ascent step to promote feasibility based on the newly updated primal variable $x_{k+1}$. Finally, the misspecified parameter $\theta_{k+1}$ is refined by performing one projected gradient step associated with the secondary VI problem (2). The steps of the algorithm are outlined in Algorithm 1.
Algorithm 1 Augmented Lagrangian Method for Misspecified VI (ALM-Misspecified VI)
1:Input: $x_{0}\in X$, $\theta_{0}\in\Theta$, $\{\rho_{k},\gamma_{k},\eta_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$.
2:Initialization: $x_{-1}\leftarrow x_{0}$, $\theta_{-1}\leftarrow\theta_{0}$, $\lambda_{0}\leftarrow\mathbf{0}$
3:for $k=0,1,2,\dots,K-1$ do
4:  $r_{k}\leftarrow F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$
5:  $x_{k+1}\leftarrow\Pi_{X}\left[x_{k}-\gamma_{k}\left(F(x_{k},\theta_{k})+r_{k}+\mathbf{J}f(x_{k},\theta_{k})[\rho_{k}f(x_{k},\theta_{k})+\lambda_{k}]_{+}\right)\right]$
6:  $\lambda_{k+1}\leftarrow\left[\lambda_{k}+\rho_{k}f(x_{k+1},\theta_{k})\right]_{+}$
7:  $\theta_{k+1}\leftarrow\Pi_{\Theta}\left[\theta_{k}-\eta_{k}H(\theta_{k})\right]$
8:end for
4 Convergence Analysis
In this section, we establish the convergence properties of Algorithm 1 under Assumption 2.1. In particular, through a step-by-step analysis, we demonstrate that our method achieves a convergence rate of $\mathcal{O}(1/K)$ in terms of the relaxed gap and infeasibility metrics. All related proofs are provided in the Appendix.
We begin by showing some preliminary bounds on the difference between the constraint enforcement term of the augmented Lagrangian function, $\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$, and the corresponding term in the standard Lagrangian, $\sum^{J}_{j=1}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$.
Lemma 4.1.
[…]
$-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$
Proof.
Proof See Appendix A for the proof. ∎
Based on the result of Lemma 4.1, we next establish a one-step bound for the iterates generated by the proposed algorithm.
Lemma 4.2 (Lipschitz continuity of $\nabla_{x}\Phi_{\rho}(\cdot,\lambda,\theta)$).
Suppose Assumption 2.1 holds. Then, for every $\rho>0$, every $\lambda\in\mathbb{R}^{J}$, and every $\theta\in\Theta$, the mapping $\nabla_{x}\Phi_{\rho}(\cdot,\lambda,\theta)$ is Lipschitz continuous on $X$. In particular, for any $x,y\in X$, $\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|\leq(\rho C_{1}+\|\lambda\|C_{2})\|x-y\|,$ where $C_{1}\triangleq\sqrt{J}\big(L_{\nabla f,x}D_{f}+L_{f,x}M_{\nabla f}\big),\quad C_{2}\triangleq\sqrt{J}\,L_{\nabla f,x},$ and $D_{f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|[f(z,\vartheta)]_{+}\|<\infty,\quad M_{\nabla f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|\nabla_{x}f(z,\vartheta)\|<\infty.$
Proof.
Proof See Appendix B for the proof. ∎
Lemma 4.3 (One-step analysis).
Let $\{(x_{k},\lambda_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 1 and suppose Assumption (2.1) holds. Then, for any $x\in X$, $\lambda\in\mathbb{R}^{J}_{+}$, the following inequality holds for any $k\geq 0$,
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)$
$\qquad+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right)+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
$\qquad+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|,$
where $r_{k}=F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$, $\gamma_{k}>0$ is the step-size, and $\rho_{k}>0$ is the penalty parameter.
Proof.
Proof See Appendix C for the proof. ∎
Note that the last term on the right-hand side of the inequality contains the product $\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$. Assuming that the learning iterates $\{\theta_{k}\}$ are converging to $\theta^{*}$, this term will be summable only if the dual iterates do not grow. In fact, in the following lemma, we establish that the dual iterates $\{\lambda_{k}\}_{k\geq 0}$ are bounded.
Lemma 4.4 (Boundedness of Dual Iterates).
Suppose Assumption 2.1 holds, and define $D_{X}\triangleq\sup_{x\in X}\|x-x^{*}\|<\infty$ and $D_{\Theta}\triangleq\sup_{\theta\in\Theta}\|\theta-\theta^{*}\|<\infty$. Consider Algorithm 1 with constant penalty $\rho_{k}\equiv\rho>0$ and primal step-size $\gamma_{k}\equiv\gamma\in\bigl(0,(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda})^{-1}\bigr)$ for some $D_{\Lambda}>0$ where $C_{1},C_{2}$ are defined in Lemma 4.2. Assume that the learning error satisfies $\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|<\infty$ and $\sum_{k=0}^{\infty}(k+1)^{2}\|\theta_{k}-\theta^{*}\|^{2}<\infty,$ and $\rho_{k}=\rho=\frac{1}{L_{\lambda\theta}^{\Phi}}$ for all $k\geq 0$. Then the dual sequence $\{\lambda_{k}\}$ is bounded.
Proof.
Proof See Appendix D for the proof. ∎
Lemma 4.4 establishes bounded dual iterates under compactness and summable learning errors when we use a constant primal step-size and a penalty parameter. We adopt this schedule henceforth, which yields a uniform bound $D_{\Lambda}\triangleq\sup_{k\geq 0}\|\lambda_{k}-\lambda^{*}\|<+\infty$. With this standing in place, we now state the main convergence result for the ergodic primal averages.
Theorem 4.5 (Convergence Rate).
Suppose Assumption 2.1 holds, and Algorithm 1 is executed under the conditions of Lemma $4.$ Then, for the ergodic average $\bar{x}_{K}=\tfrac{1}{K}\sum_{k=1}^{K}x_{k}$, the following hold:
  • (i)
    (Infeasibility:) $\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}\leq\frac{C_{\mathrm{feas}}}{K}$.
  • (ii)
    (Relaxed gap:) $\widetilde{\mathrm{Gap}}(\bar{x}_{K},\theta^{*})\triangleq\sup_{x\in\mathcal{X}_{\epsilon_{K}}(\theta^{*})}F(x,\theta^{*})^{\top}(\bar{x}_{K}-x)\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}+D_{\Lambda}\frac{C_{\mathrm{feas}}}{K}+\frac{\rho}{2}\left(\frac{C_{\mathrm{feas}}}{K}\right)^{2},$
where $C^{\prime}_{T}\triangleq\tfrac{1}{2\gamma}D_{X}^{2}+2L_{F,\theta}D_{X}D_{\Theta}$, $C^{\prime}_{R}(x,0)=\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}$, and $C_{\mathrm{feas}}=C^{\prime}_{T}+\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J+\|\lambda_{0}-\lambda^{*}\|^{2}}{\rho}$.
Proof.
Proof See Appendix E for the proof. ∎
Remark 4.1.
Based on Assumption 2.1-(ii), the secondary VI problem (2) has a Lipschitz continuous and strongly monotone operator which implies that the sequence $\{\theta_{k}\}_{k\geq 0}$ generated by the projected gradient step in line 7 of Algorithm 1 converges to the optimal unique parameter $\theta^{*}$ at a linear convergence rate (e.g. see [7, Proposition 26.16]); hence, the learning error accumulations $\sum_{k}\|\theta_{k}-\theta^{*}\|$ and $\sum_{k}(k+2)^{2}\|\theta_{k}-\theta^{*}\|^{2}$ in the upper bounds of the result of Theorem 4.5 are finite with a uniform bound. Therefore, the result of Theorem 4.5 implies that achieving $\epsilon$-infeasibility and $\epsilon$-relaxed gap requires at most running $\mathcal{O}(1/\epsilon)$ iterations of Algorithm 1. We finally remark that the update of misspecified parameter $\theta$ in the algorithm can be replaced with other methods, such as extra-gradient method, as long as $\theta_{k}\to\theta^{*}$ at a linear rate.
5 Numerical Experiment
In this section, we test the performance of our proposed method by solving the Cournot model described in Section 1.3 and compare with the Tikhonov regularized VI scheme [9, 3] and Extragradient method by [3].
We consider a market with $N$ firms, where each firm $i\in\{1,\dots,N\}$ produces $D$ homogeneous products so the total decision variable dimension is $n=ND$. For each product $d=1,\dots,D$ of firm $i$, the capacity is $\text{Cap}_{i}=5$ and the quadratic cost function $c_{i}(x_{i})=\tfrac{1}{2}r_{i}x_{i}^{2}+g_{i}x_{i}$ where $r_{i},g_{i}$ are randomly generated from uniform distributions $U[1,10]$ and $U[5,20]$, respectively. Moreover, $p(X;b^{*})=a^{*}-b^{*}X$ where $X=\sum_{i=1}^{N}x_{i}$, we let $a^{*}=100$, and $b^{*}$ is the solution of the least-square minimization (3) with synthetic observations generated from a true model $X_{t}\sim U[2,20]$, $p_{\mathrm{obs},t}=a^{*}-b^{*}X_{t}$, for $t=1,\dots,T=300$. All the firms should also satisfy the coupling constraint of the form $f(\mathbf{x},b^{*})\triangleq p(X;b^{*})-\delta\leq 0$ for some $\delta>0$.
We implemented our proposed ALM-Misspecified VI method and compare it with the misspecified Tikhonov regularized VI [9, 3] and Extragradient (EG) method by [3]. For all methods, the misspecified parameter is updated using one projected gradient step at each iteration. To handle the coupling constraint for the other two methods, we employ the Lagrangian reformulation for each player and construct the corresponding VI operator by concatenating all the players’ gradient vectors of the Lagrangian function with respect to primal and dual variables. We consider three problem instances with $(N,D)\in\{(50,5),(50,10),(100,10)\}$ and run each method for a sufficiently large number of iterations with constant step-sizes chosen according to the theoretical findings. Figure 1, shows the performance of the methods in terms of the VI gap $\widetilde{\mathrm{Gap}}(\bar{x}_{K},b^{*})$ and infeasibility $\mathbf{1}^{\top}[f(\bar{x}_{K},b^{*})]_{+}$. Our proposed method (ALM-Misspecified VI) exhibits a superior performance across all instances, while the Lagrangian-Tikhonov and EG-Lagrangian baselines converge more slowly and deteriorate more noticeably as $(N,D)$ grow.
Figure 1: Cournot instances with $(N,D)\in\{(50,5),(50,10),(100,10)\}$.
The plots show the VI-Gap (left) and infeasibility (right) for ALM-Misspecified VI (solid), Lagrangian–Tikhonov (dashed with circles), and EG–Lagrangian (dotted with crosses).
Appendix A Proof of Lemma 4.1
Proof.
Proof Let $j$ be an index in $\{1,\dots,J\}$. For any $j^{th}$ coordinate using Algorithm 1 (line 6), the dual update is $\lambda_{k+1}^{(j)}=\Bigl[\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})\Bigr]_{+},$ where $\cdot$$_{+}$ denotes the projection onto $\mathbb{R}_{+}$. We now consider two cases, when $f_{j}^{\prime}s$ are active, we have $\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})>0.$ Then the projection acts as the identity, so we have
$\lambda_{k+1}^{(j)}$ $=\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})\quad\implies\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}f_{j}(x_{k+1},\theta_{k}).$
Recall the definition 2.1, then we have
$\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}=\begin{cases}f_{j}(x_{k+1},\theta_{k})&\text{if }\rho_{k}f_{j}(x_{k+1},\theta_{k})+\lambda_{k}^{(j)}\geq 0\\ -\dfrac{\lambda_{k}^{(j)}}{\rho_{k}}&\text{if }\rho_{k}f_{j}(x_{k+1},\theta_{k})+\lambda_{k}^{(j)}<0.\end{cases}$ (6)
Thus, we conclude that in the active case $\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}.$ Again, if $f_{j}(x_{k+1},\theta_{k})^{\prime}s$ are inactive, then we have $\lambda_{k}^{(j)}+\rho_{k}f_{j}(x_{k+1},\theta_{k})<0$, hence, $f_{j}(x_{k+1},\theta_{k})<-\frac{\lambda_{k}^{(j)}}{\rho_{k}}.$ Thus, we have $\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}=-\lambda_{k}^{(j)}.$ Therefore, in inactive case also we have $\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}=\rho_{k}\Bigl[\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\Bigr]_{j}.$ Since the above equality holds for both active and inactive cases, we can write in full vector notation that $\lambda_{k+1}-\lambda_{k}=\rho_{k}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k}).$ Now by adding and subtracting $\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$, we get
$\lambda_{k+1}-\lambda_{k}=\rho_{k}\Bigl[e_{k}+\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)\Bigr],$ (7)
where in the last equality, we define $e_{k}=\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)-\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$. For any $\lambda\in\mathbb{R}_{+}^{J}$, using the three-point identity and dividing both sides by $\rho_{k}$ we get
$\frac{1}{\rho_{k}}(\lambda_{k+1}-\lambda)^{\top}(\lambda_{k+1}-\lambda_{k})$ $=\frac{1}{2\rho_{k}}\Bigl[\|\lambda_{k+1}-\lambda\|^{2}-\|\lambda_{k}-\lambda\|^{2}+\|\lambda_{k+1}-\lambda_{k}\|^{2}\Bigr].$ (8)
Using the expression in (7) for $\lambda_{k+1}-\lambda_{k}$, we see
$\frac{1}{\rho_{k}}(\lambda_{k+1}-\lambda)^{\top}(\lambda_{k+1}-\lambda_{k})$ $=(\lambda_{k+1}-\lambda)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+(\lambda_{k+1}-\lambda)^{\top}e_{k}.$ (9)
Notice that, using (8), we can rewrite (9) as follows
$\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}$ $=\frac{1}{2\rho_{k}}\|\lambda_{k}-\lambda\|^{2}-\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda\|^{2}+(\lambda_{k+1}-\lambda)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+(\lambda_{k+1}-\lambda)^{\top}e_{k}$
$\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)+(\lambda_{k+1}-\lambda)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$
$\quad+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|,$ (10)
[…]
$\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}$ $=\rho_{k}f_{j}(x_{k+1},\theta_{k})=\rho_{k}\left[f_{j}(x_{k+1},\theta^{*})+\left(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\right)\right].$
Since $a=b+c$, squaring both sides and using the fact that $(b+c)^{2}\geq\frac{1}{2}b^{2}-c^{2}$, we obtain
$\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}$ $=\rho_{k}^{2}\left(f_{j}(x_{k+1},\theta_{k})\right)^{2}$
$\geq\rho_{k}^{2}\left(\frac{1}{2}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-\left(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\right)^{2}\right).$
Finally, invoking Lipschitz continuity of $f_{j}(x,\theta)$ in $\theta$ from Assumption 2.1(iii), we have
$\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}\geq\rho_{k}^{2}\left[\frac{1}{2}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right].$
Summing over $j\in J_{k}^{+}$ and multiplying by $1/(2\rho_{k})$, we have
$\frac{1}{2\rho_{k}}\sum_{j\in J_{k}^{+}}\left(\lambda_{k+1}^{(j)}-\lambda_{k}^{(j)}\right)^{2}\geq\frac{\rho_{k}}{4}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}-\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}.$
If $j\in J_{k}^{-}$, then $\lambda^{(j)}_{k+1}=0\implies\left(\lambda^{(j)}_{k+1}-\lambda^{(j)}_{k}\right)^{2}=\left(\lambda^{(j)}_{k}\right)^{2}$. Multiplying both sides by $\frac{1}{2\rho_{k}}$, and combining both of the above inequalities, we get
$\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}\geq\tfrac{\rho_{k}}{4}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}+\tfrac{1}{2\rho_{k}}\sum_{j\in J_{k}^{-}}(\lambda_{k}^{(j)})^{2}-\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}.$
After rearranging, this gives us
$-\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right)\leq-\Biggl[\tfrac{\rho_{k}}{4}\sum_{j\in J_{k}^{+}}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}+\tfrac{1}{2\rho_{k}}\sum_{j\in J_{k}^{-}}(\lambda_{k}^{(j)})^{2}\Biggr].$ (11)
Now, notice that
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)$ $=\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)+\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)$
$=\sum_{j\in J_{k}^{+}}\Bigl(\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta_{k})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta_{k})\Bigr)-\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}$
$\qquad+\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr).$
For each $j\in J_{k}^{+}$, we add and subtract the term $\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})$ inside the summation, hence we obtain
$\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$ $=\sum_{j\in J_{k}^{+}}\Bigl(\tfrac{\rho_{k}}{2}\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})\Bigr)-\sum_{j\in J_{k}^{-}}\frac{(\lambda_{k}^{(j)})^{2}}{2\rho_{k}}+E_{k},$ (12)
where $E_{k}\triangleq\sum_{j\in J_{k}^{+}}\left(\tfrac{\rho_{k}}{2}\bigl(\left(f_{j}(x_{k+1},\theta_{k})\right)^{2}-\left(f_{j}(x_{k+1},\theta^{*})\right)^{2}\bigr)+\lambda_{k}^{(j)}\bigl(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\bigr)\right)+\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta_{k}\bigr)$. Now using the facts from (6) and (12), we can expand as follows
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad=\sum_{j\in J_{k}^{+}}\Bigl(\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})\Bigr)-\sum_{j\in J_{k}^{-}}\tfrac{(\lambda_{k}^{(j)})^{2}}{2\rho_{k}}-\sum_{j\in J_{k}^{+}}\bigl(\lambda^{(j)}_{k+1}-\lambda^{(j)}\bigr)f_{j}(x_{k+1},\theta^{*})$
$\qquad-\sum_{j\in J_{k}^{-}}\bigl(\lambda^{(j)}_{k+1}-\lambda^{(j)}\bigr)\Bigl(-\tfrac{\lambda^{(j)}_{k}}{\rho_{k}}\Bigr)-\sum_{j\in J_{k}^{+}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\sum_{j\in J_{k}^{-}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+E_{k}.$
Next, by grouping the positive and negative index sets on the RHS separately, we get
$=\sum_{j\in J_{k}^{+}}\Bigl(\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})\Bigr)-\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}-\sum_{j\in J_{k}^{-}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$
$\quad-\sum_{j\in J_{k}^{+}}\left(\lambda^{(j)}_{k}+\rho_{k}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}\right)f_{j}(x_{k+1},\theta^{*})-\sum_{j\in J_{k}^{-}}\left(-\lambda^{(j)}\right)\left(-\frac{\lambda^{(j)}_{k}}{\rho_{k}}\right)-\sum_{j\in J^{+}_{k}}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+E_{k}.$
Now, observe that the terms inside the sum over $J_{k}^{+}$ can be combined, since both involve $f_{j}(x_{k+1},\theta^{*})$. After cancellation we arrive at
$=\sum_{j\in J_{k}^{+}}\Bigl[\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}+\lambda_{k}^{(j)}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$
$\quad-\left(\lambda^{(j)}_{k}+\rho_{k}f_{j}(x_{k+1},\theta^{*})-\lambda^{(j)}\right)f_{j}(x_{k+1},\theta^{*})\Bigr]-\sum_{j\in J_{k}^{-}}\left(\tfrac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}+\lambda^{(j)}\Bigl(f_{j}(x_{k+1},\theta^{*})+\tfrac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\right)+E_{k}.$
Finally, simplifying the quadratic terms in $f_{j}(x_{k+1},\theta^{*})$, we can conclude that
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad=-\sum_{j\in J_{k}^{+}}\tfrac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*})\bigr)^{2}-\sum_{j\in J_{k}^{-}}\left(\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}+\lambda^{(j)}\Bigl(f_{j}(x_{k+1},\theta^{*})+\frac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\right)+E_{k}.$ (13)
Note that, for any $\lambda^{(j)}\geq 0$ and $\forall j\in J_{k}^{-}$, we have $\Bigl(f_{j}(x_{k+1},\theta^{*})+\frac{\lambda_{k}^{(j)}}{\rho_{k}}\Bigr)\leq 0$. Then from (A), we have
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad\geq-\left(\sum_{j\in J_{k}^{+}}\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*}\bigr)^{2}+\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}\right)+E_{k}$
$\quad\geq-\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right)+E_{k},$ (14)
where the last inequality follows from (11). Therefore, we have
- $\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad\leq\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right)-E_{k}.$ (15)
Recall that $E_{k}=\sum_{j\in J_{k}^{+}}\left(\frac{\rho_{k}}{2}\Bigl(f_{j}(x_{k+1},\theta_{k})^{2}-f_{j}(x_{k+1},\theta^{*})^{2}\Bigr)+\lambda_{k}^{(j)}\Bigl(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\Bigr)\right)+\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k}).$ Using the triangle inequality and $a^{2}-b^{2}=(a-b)(a+b)$, we obtain
$|E_{k}|$ $\leq\sum_{j\in J_{k}^{+}}\frac{\rho_{k}}{2}|f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})||f_{j}(x_{k+1},\theta_{k})+f_{j}(x_{k+1},\theta^{*})|$
$\qquad+\sum_{j\in J_{k}^{+}}|\lambda_{k}^{(j)}||f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})|+\left|\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\right|.$ (16)
Recall that $\Phi_{\rho}(x,\lambda,\theta)=\sum_{j=1}^{J}\phi_{\rho}(f_{j}(x,\theta),\lambda_{j})$, moreover, one can observe that for any $u_{j},u_{j}^{\prime}\in\mathbb{R}$, $|\phi_{\rho}(u_{j},v_{j})-\phi_{\rho}(u_{j}^{\prime},v_{j})|\leq(|v_{j}|+\rho D_{f})|u_{j}-u_{j}^{\prime}|$. Therefore, by Lipschitz continuity of $f_{j}(x,\theta)$ in $\theta$ we conclude that for any $\theta,\theta^{\prime}\in\Theta$, $|\phi_{\rho}(f_{j}(x,\theta),\lambda^{(j)})-\phi_{\rho}(f_{j}(x,\theta^{\prime}),\lambda^{(j)})|\leq(|\lambda^{(j)}|+\rho D_{f})L_{f,\theta}\|\theta-\theta^{\prime}\|$ which summing over $j$ implies that
$|\Phi_{\rho}(x,\lambda,\theta)-\Phi_{\rho}(x,\lambda,\theta^{\prime})|\leq L_{f,\theta}(J\rho D_{f}+\sqrt{J}\|\lambda\|)\|\theta-\theta^{\prime}\|.$ (17)
Using the above inequality within (A) and by Lipschitz continuity of $f_{j}(x,\theta)$ in $\theta$ and the bound $|f_{j}(x,\theta)|\leq D_{f}$, we conclude that
$|E_{k}|\leq 2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (18)
Now using (A) and (18) in (A), we can show
$-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$ $\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
$\quad+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (19)
Appendix B Proof of Lemma 4.2
Proof.
Proof Recall from Definition 2.1 and (4), for any $x\in X$, we have
$\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)=\sum_{j=1}^{J}[\rho f_{j}(x,\theta)+\lambda^{(j)}]_{+}\,\nabla_{x}f_{j}(x,\theta)=\nabla_{x}f(x,\theta)^{\top}[\rho f(x,\theta)+\lambda]_{+}.$
Fix $\rho>0$, $\lambda\in\mathbb{R}^{J}$, and $\theta\in\Theta$. Then for any $x,y\in X$, we have
$\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)$ $=\nabla_{x}f(x,\theta)^{\top}[\rho f(x,\theta)+\lambda]_{+}-\nabla_{x}f(y,\theta)^{\top}[\rho f(y,\theta)+\lambda]_{+}$
$=\big(\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\big)^{\top}[\rho f(x,\theta)+\lambda]_{+}$
$\quad+\nabla_{x}f(y,\theta)^{\top}\Big([\rho f(x,\theta)+\lambda]_{+}-[\rho f(y,\theta)+\lambda]_{+}\Big),$
where we added and subtracted $\nabla_{x}f(y,\theta)^{\top}[\rho f(x,\theta)+\lambda]_{+}$. Taking norms and applying the triangle inequality yields
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|$ $\leq\|\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\|\,\|[\rho f(x,\theta)+\lambda]_{+}\|$
$\quad+\|\nabla_{x}f(y,\theta)\|\,\|[\rho f(x,\theta)+\lambda]_{+}-[\rho f(y,\theta)+\lambda]_{+}\|.$ (20)
Since $\cdot$$_{+}$ is the projection onto $\mathbb{R}^{J}_{+}$, it is non-expansive. Therefore,
$\|f(x,\theta)-f(y,\theta)\|=\Big(\sum_{j=1}^{J}|f_{j}(x,\theta)-f_{j}(y,\theta)|^{2}\Big)^{1/2}\leq\sqrt{J}\,L_{f,x}\|x-y\|.$
Consequently, it follows that
$\|[\rho f(x,\theta)+\lambda]_{+}-[\rho f(y,\theta)+\lambda]_{+}\|\leq\rho\sqrt{J}\,L_{f,x}\|x-y\|.$
Similarly, by Assumption 2.1(iii), each $\nabla_{x}f_{j}(\cdot,\theta)$ is Lipschitz continuous in $x$ with constant $L_{\nabla f,x}$. Therefore, the Jacobian matrix satisfies
$\|\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\|\leq\sqrt{J}\,L_{\nabla f,x}\|x-y\|.$
Substituting these bounds into (B), we obtain
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|$ $\leq\sqrt{J}\,L_{\nabla f,x}\|x-y\|\,\|[\rho f(x,\theta)+\lambda]_{+}\|$
$\quad+\rho\sqrt{J}\,L_{f,x}\|x-y\|\,\|\nabla_{x}f(y,\theta)\|.$
Now, since $X\times\Theta$ is compact and the mappings $(z,\theta)\mapsto f(z,\theta)$ and $(z,\theta)\mapsto\nabla_{x}f(z,\theta)$ are continuous, the quantities $D_{f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|[f(z,\vartheta)]_{+}\|<\infty,\quad M_{\nabla f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|\nabla_{x}f(z,\vartheta)\|<\infty,\$are well defined and finite. Moreover, $\|[\rho f(x,\theta)+\lambda]_{+}\|\leq\rho\|[f(x,\theta)]_{+}\|+\|\lambda\|\leq\rho D_{f}+\|\lambda\|,$ and $\|\nabla_{x}f(y,\theta)\|\leq M_{\nabla f}$. Hence, we get
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|$ $\leq\sqrt{J}\,L_{\nabla f,x}(\rho D_{f}+\|\lambda\|)\|x-y\|+\rho\sqrt{J}\,L_{f,x}M_{\nabla f}\|x-y\|.$
Rearranging terms gives
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|\leq\Big(\rho C_{1}+\|\lambda\|C_{2}\Big)\|x-y\|,$
where $C_{1}\triangleq\sqrt{J}\big(L_{\nabla f,x}D_{f}+L_{f,x}M_{\nabla f}\big),\quad C_{2}\triangleq\sqrt{J}\,L_{\nabla f,x}.$ Thus, $\nabla_{x}\Phi_{\rho}(\cdot,\lambda,\theta)$ is Lipschitz continuous on $X$. ∎
Appendix C Proof of Lemma 4.3
Proof.
Proof Let $x\in X$ and $\lambda\in\mathbb{R}^{J}_{+}$ be arbitrary vectors. From Algorithm 1 (line 5) we have:
$(x_{k+1}-x)^{\top}\Bigg(x_{k+1}-x_{k}+\gamma_{k}\Big(F(x_{k},\theta_{k})+r_{k}+\mathbf{J}f(x_{k},\theta_{k})\Big[\rho_{k}f(x_{k},\theta_{k})+\lambda_{k}\Big]_{+}\Big)\Bigg)\leq 0.$ (21)
Considering (21), to find a lower bound for $\left(x_{k+1}-x\right)^{\top}\Big(F\left(x_{k},\theta_{k}\right)+r_{k}\Big)$, we add and subtract $(x_{k+1}-x)^{\top}F(x_{k+1},\theta_{k+1})$, obtaining:
$\left(x_{k+1}-x\right)^{\top}\Big(F\left(x_{k},\theta_{k}\right)+r_{k}\Big)\pm(x_{k+1}-x)^{\top}F(x_{k+1},\theta_{k+1})$
$\quad=\langle r_{k},x_{k}-x\rangle-\langle r_{k+1},x_{k+1}-x\rangle+\langle r_{k},x_{k+1}-x_{k}\rangle+\langle F(x_{k+1},\theta_{k+1}),x_{k+1}-x\rangle$
$\quad=\left(x_{k+1}-x\right)^{\top}\Big(F(x_{k+1},\theta^{*})+[F(x_{k+1},\theta_{k+1})-F(x_{k+1},\theta^{*})]\Big)+\langle r_{k},x_{k}-x\rangle$
$\qquad-\langle r_{k+1},x_{k+1}-x\rangle+\langle r_{k},x_{k+1}-x_{k}\rangle$
$\quad\geq\left(x_{k+1}-x\right)^{\top}F(x,\theta^{*})+\left(x_{k+1}-x\right)^{\top}[F(x_{k+1},\theta_{k+1})-F(x_{k+1},\theta^{*})]+\langle r_{k},x_{k}-x\rangle$
$\qquad-\langle r_{k+1},x_{k+1}-x\rangle+\langle r_{k},x_{k+1}-x_{k}\rangle,$
where in the last inequality, we use the monotonicity of $F(\cdot,\theta^{*})$. Next, using Cauchy-Schwarz inequality and Lipschitz continuity of $F(x,\cdot)$ with respect to $\theta$, we can show:
$\left(x_{k+1}-x\right)^{\top}\left(F(x_{k},\theta_{k})+r_{k}\right)$
$\quad\geq\left(x_{k+1}-x\right)^{\top}F(x,\theta^{*})-\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)\|x_{k+1}-x_{k}\|$
$\qquad-L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+\langle r_{k},x_{k}-x\rangle-\langle r_{k+1},x_{k+1}-x\rangle$ (22)
Taking into account 4, by adding and subtracting $\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})$, we can obtain the following.
$(x_{k+1}-x)^{\top}\mathbf{J}f(x_{k},\theta_{k})\Big[\rho_{k}f(x_{k},\theta_{k})+\lambda_{k}\Big]_{+}$
$\quad=(x_{k+1}-x_{k})^{\top}\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k})+(x_{k}-x)^{\top}\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k})$
$\quad=(x_{k+1}-x_{k})^{\top}\Big(\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})+[\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k})-\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})]\Big)$
$\qquad+(x_{k}-x)^{\top}\Big(\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})+[\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k})-\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})]\Big)$
$\quad\geq(x_{k+1}-x_{k})^{\top}\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})+(x_{k+1}-x_{k})^{\top}[\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k})-\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})]$
$\qquad+\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})+(x_{k}-x)^{\top}[\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k})-\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})],$
where in the last inequality, we use the convexity of $\Phi_{\rho_{k}}(\cdot,\lambda_{k},\theta^{*})$. Next, using the Cauchy-Schwarz inequality and Lipschitz continuity of $\nabla_{x}\Phi_{\rho_{k}}(x,\lambda,\cdot)$ with respect $\theta$ with a constant $L^{\Phi}_{x\theta}$, we can show
$(x_{k+1}-x)^{\top}\mathbf{J}f(x_{k},\theta_{k})\Big[\rho_{k}f(x_{k},\theta_{k})+\lambda_{k}\Big]_{+}\geq(x_{k+1}-x_{k})^{\top}\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})$
$\quad+\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})-L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big).$ (23)
Moreover, using the three-point equality, we can write:
$(x_{k+1}-x)^{\top}(x_{k+1}-x_{k})=\frac{1}{2}\Big(\|x_{k+1}-x\|^{2}-\|x_{k}-x\|^{2}+\|x_{k+1}-x_{k}\|^{2}\Big).$ (24)
Using (C), (C), and (24) in (21), we have:
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}-\|x_{k+1}-x_{k}\|^{2}\Big)-(x_{k+1}-x_{k})^{\top}\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})$
$\qquad+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)$
$\qquad+\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle.$ (25)
Next, using the Lemma 4.2, we have $\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})\leq\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})+\langle\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*}),x_{k+1}-x_{k}\rangle+\frac{\rho_{k}C_{1}+\|\lambda_{k}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}$. Substituting this estimate into (C) yields
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}-\|x_{k+1}-x_{k}\|^{2}\Big)+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle$
$\qquad+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)$
$\qquad+\frac{\rho_{k}C_{1}+\|\lambda_{k}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big).$ (26)
To treat the term involving $\|\lambda_{k}\|$, we first add and subtract $\lambda^{*}$ inside the norm and use the triangle inequality to obtain
$\frac{\rho_{k}C_{1}+\|\lambda_{k}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}$ $=\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}+(\lambda_{k}-\lambda^{*})\|C_{2}}{2}\right)\|x_{k+1}-x_{k}\|^{2}$
$\leq\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}.$ (27)
Using (A) from Lemma (4.1) and substituting (C) into (C), we obtain
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}-\|x_{k+1}-x_{k}\|^{2}\Big)+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
$\qquad+\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)$
$\qquad+\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (28)
Note that applying Young’s inequality to the term $\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)$, we have
$\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)$
$\quad\leq\frac{L_{F,x}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,x}}{2}\|x_{k}-x_{k-1}\|^{2}+\frac{L_{F,\theta}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
$\quad=\left(\frac{L_{F,x}+L_{F,\theta}}{2}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,x}}{2}\|x_{k}-x_{k-1}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}.$
Now let us add and subtract $\frac{L_{F,x}}{2}\|x_{k+1}-x_{k}\|^{2}$ and $\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$ with the respective terms to obtain
$\left(\frac{L_{F,x}+L_{F,\theta}}{2}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,x}}{2}\|x_{k}-x_{k-1}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
$\quad=\left(\frac{2L_{F,x}+L_{F,\theta}}{2}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right).$
Putting all terms together in (C), we get
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)$
$\qquad+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right)$
$\qquad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (29)
Appendix D Proof of Lemma 4.4
Proof.
Proof From (C), setting $(x,\lambda)=(x^{*},\lambda^{*})$ gives, for all $k\geq 0$
$0\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda^{*}\|^{2}-\|\lambda_{k+1}-\lambda^{*}\|^{2}\Big)+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}-\langle r_{k},x_{k}-x^{*}\rangle$
$\qquad+\langle r_{k+1},x_{k+1}-x^{*}\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\Big)$
$\qquad+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right)$
$\qquad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (30)
To bound the cross term $L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|$, we apply Young’s inequality for any $\alpha_{k}>0$
$L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|\leq\frac{\alpha_{k}L^{\Phi}_{\lambda\theta}}{2}\|\lambda_{k+1}-\lambda^{*}\|^{2}+\frac{L^{\Phi}_{\lambda\theta}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}.$
Using the above inequality in (D), multiplying both sides by $2\rho_{k}$ and rearranging the terms, we obtain
$\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq\|\lambda_{k}-\lambda^{*}\|^{2}+\rho_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\rho_{k}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}+2\rho_{k}\left(T_{k}+R_{k}\right),$ (31)
where
$T_{k}$ $\triangleq\tfrac{1}{2\gamma_{k}}\left(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\right)+\tfrac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)+\langle r_{k+1},x_{k+1}-x^{*}\rangle-\langle r_{k},x_{k}-x^{*}\rangle,$
$R_{k}$ $\triangleq L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\big)+\tfrac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
$\quad+\tfrac{L^{\Phi}_{\lambda\theta}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$
Assume now that $0<\rho_{k}L_{\lambda\theta}^{\Phi}\alpha_{k}<1$. Then, dividing both sides of (D) by $\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)$ we get
$\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq\frac{1}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\|\lambda_{k}-\lambda^{*}\|^{2}+\frac{\rho_{k}}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\frac{\rho_{k}}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\frac{2\rho_{k}}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\left(T_{k}+R_{k}\right).$ (32)
We now provide upper bounds the auxiliary terms $T_{k}$ and $R_{k}$. For $T_{k}$ summing from $k=0,\dots,K-1$, using $x_{-1}=x_{0}$ and $r_{0}=0$, and $\gamma_{k}=\gamma$ we get
$\sum_{k=0}^{K-1}T_{k}$ $=\frac{1}{2\gamma}\Bigl(\|x_{0}-x^{*}\|^{2}-\|x_{K}-x^{*}\|^{2}\Bigr)+\frac{L_{F,x}}{2}\bigl(\|x_{0}-x_{-1}\|^{2}-\|x_{K}-x_{K-1}\|^{2}\bigr)$
$\quad+\sum_{k=0}^{K-1}\Bigl(\Bigl\langle r_{k+1},x_{k+1}-x^{*}\Bigr\rangle-\Bigl\langle r_{k},x_{k}-x^{*}\Bigr\rangle\Bigr)$
$=\frac{1}{2\gamma}\Bigl(\|x_{0}-x^{*}\|^{2}-\|x_{K}-x^{*}\|^{2}\Bigr)+\frac{L_{F,x}}{2}\bigl(\|x_{0}-x_{-1}\|^{2}-\|x_{K}-x_{K-1}\|^{2}\bigr)+\Bigl\langle r_{K},x_{K}-x^{*}\Bigr\rangle.$ (33)
By Lipschitzness of $F$ in $(x,\theta)$ and recalling that $r_{K}=F(x_{K},\theta_{K})-F(x_{K-1},\theta_{K-1})$, one can easily show that $\|r_{K}\|\leq L_{F,x}\|x_{K}-x_{K-1}\|+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|.$ Hence, by Cauchy–Schwarz and Young’s inequality
$\langle r_{K},x_{K}-x^{*}\rangle$ $\leq\left(L_{F,x}\|x_{K}-x_{K-1}\|+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|\right)\|x_{K}-x^{*}\|$
$\leq\tfrac{L_{F,x}}{2}\|x_{K}-x_{K-1}\|^{2}+\tfrac{L_{F,x}}{2}\|x_{K}-x^{*}\|^{2}+\tfrac{L_{F,\theta}}{2}\|\theta_{K}-\theta_{K-1}\|^{2}+\tfrac{L_{F,\theta}}{2}\|x_{K}-x^{*}\|^{2}$
$=\tfrac{L_{F,x}}{2}\|x_{K}-x_{K-1}\|^{2}+\tfrac{L_{F,\theta}}{2}\|\theta_{K}-\theta_{K-1}\|^{2}+\tfrac{L_{F,x}+L_{F,\theta}}{2}\|x_{K}-x^{*}\|^{2}.$
From the definition of $T_{k}$ and telescoping with $x_{-1}=x_{0}$ and $r_{0}=0$ in (D), we have
$\sum_{k=0}^{K-1}T_{k}$ $\leq\tfrac{1}{2\gamma}\|x_{0}-x^{*}\|^{2}+\left(\tfrac{L_{F,x}+L_{F,\theta}}{2}-\tfrac{1}{2\gamma}\right)\|x_{K}-x^{*}\|^{2}+\tfrac{L_{F,\theta}}{2}\|\theta_{K}-\theta_{K-1}\|^{2}.$ (34)
Choosing $\gamma\in\Big(0,(L_{F,x}+L_{F,\theta})^{-1}\Big)$ the coefficient of $\|x_{K}-x^{*}\|^{2}$ on the right-hand side of (34) is non-positive and may be discarded. Hence, $\sum_{k=0}^{K-1}T_{k}\leq\frac{1}{2\gamma}\|x_{0}-x^{*}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{K}-\theta_{K-1}\|^{2}.$ By the compactness of $\Theta$, $\|\theta_{K}-\theta_{K-1}\|\leq\|\theta_{K}-\theta^{*}\|+\|\theta_{K-1}-\theta^{*}\|\leq 2D_{\Theta}$, it follows that
$\sum_{k=0}^{K-1}T_{k}\leq\frac{1}{2\gamma}\|x_{0}-x^{*}\|^{2}+2L_{F,\theta}D_{\Theta}^{2}=:C_{T},$
where $C_{T}<\infty$ is independent of $K$. Therefore, for every $K\geq 1$, $\sum_{k=0}^{K-1}T_{k}\leq C_{T}.$
Next, we estimate $R_{k}$. Again, by the compactness of $X$, $\|x_{k+1}-x^{*}\|\leq D_{x}$ and $\|x_{k+1}-x_{k}\|\leq\|x_{k+1}-x^{*}\|+\|x_{k}-x^{*}\|\leq 2D_{x}$. Therefore, for every $k\geq 0$,
$R_{k}$ $\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+L_{x\theta}^{\Phi}\|\theta_{k}-\theta^{*}\|\big(2D_{X}+D_{X}\big)+\frac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
$\quad+\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J\rho_{k}}{2}L_{f,\theta}^{2}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}\big(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|\big)\|\theta_{k}-\theta^{*}\|$
$\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho_{k}D_{f}L_{f,\theta}\Big)\|\theta_{k}-\theta^{*}\|+\frac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
$\quad+\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho_{k}}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}+2\sqrt{J}L_{f,\theta}\|\lambda_{k}\|\|\theta_{k}-\theta^{*}\|.$ (35)
Using $\|\theta_{k}-\theta_{k-1}\|^{2}\leq 2\|\theta_{k}-\theta^{*}\|^{2}+2\|\theta_{k-1}-\theta^{*}\|^{2},$ we further obtain
$R_{k}$ $\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho_{k}D_{f}L_{f,\theta}\Big)\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\big(\|\theta_{k}-\theta^{*}\|^{2}+\|\theta_{k-1}-\theta^{*}\|^{2}\big)+\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho_{k}}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}$
$\quad+2\sqrt{J}L_{f,\theta}\|\lambda_{k}\|\,\|\theta_{k}-\theta^{*}\|.$ (36)
From the induction hypothesis, we have that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$, $k=0,1,\dots,K-1$, hence, by the triangle inequality, $\|\lambda_{k}\|\leq\|\lambda_{k}-\lambda^{*}\|+\|\lambda^{*}\|\leq D_{\Lambda}+\|\lambda^{*}\|$, $k=0,1,\dots,K-1.$ Therefore, for $k=0,1,\dots,K-1$, (D) implies
$R_{k}$ $\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho D_{f}L_{f,\theta}+2\sqrt{J}L_{f,\theta}(D_{\Lambda}+\|\lambda^{*}\|)\Big)\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\big(\|\theta_{k}-\theta^{*}\|^{2}+\|\theta_{k-1}-\theta^{*}\|^{2}\big)+\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}.$ (37)
Summing from $k=0$ to (K1)(K-1) and noting that $\theta_{-1}=\theta_{0}$, we obtain
$\sum_{k=0}^{K-1}R_{k}$ $\leq L_{F,\theta}D_{X}\sum_{k=1}^{K}\|\theta_{k}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho D_{f}L_{f,\theta}+2\sqrt{J}L_{f,\theta}(D_{\Lambda}+\|\lambda^{*}\|)\Big)\sum_{k=0}^{K-1}\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\Big(\|\theta_{0}-\theta^{*}\|^{2}+2\sum_{k=0}^{K-1}\|\theta_{k}-\theta^{*}\|^{2}\Big)+\sum_{k=0}^{K-1}\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}=:C_{R}(D_{\Lambda}).$
Since $\alpha_{k}=\frac{1}{(k+2)^{2}}\in(0,1)$, by the learning–error summability assumption, the right-hand side is finite for each $K$, and thus $\sum_{k=0}^{K-1}R_{k}\leq C_{R}(D_{\Lambda}).$
Now define $\delta_{k}=\frac{1}{1-\rho_{k}L_{\lambda\theta}^{\Phi}\alpha_{k}}$, then (D), becomes
$\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq\delta_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+\delta_{k}\rho_{k}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\delta_{k}\rho_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma}\right)\|x_{k+1}-x_{k}\|^{2}+2\delta_{k}\rho_{k}(T_{k}+R_{k}).$ (38)
Next, define the sequence $\{t_{k}\}_{k\geq 0}$ such that $t_{0}=1$ and $t_{k+1}=\frac{t_{k}}{\delta_{k}}\text{ for }k\geq 0$. Based on the parameter selections $\rho_{k}=\rho=\frac{1}{L_{\lambda\theta}^{\Phi}}$ and $\alpha_{k}=\frac{1}{(k+2)^{2}}$, we conclude that $t_{k}=\frac{t_{0}(k+2)}{2(k+1)}$ which implies that $t_{k}\in[\frac{1}{2},1]$ for any $k\geq 0$. Multiplying both sides of (D) by $t_{k+1}$, we obtain
$t_{k+1}\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq t_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+\rho_{k}t_{k}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\rho_{k}t_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma}\right)\|x_{k+1}-x_{k}\|^{2}+2\rho_{k}t_{k}(T_{k}+R_{k}).$ (39)
We prove the boundedness of the dual iterates by induction, showing that there exists $D_{\Lambda}>0$ such that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$ for all $k\geq 0.$ Assume, as the induction hypothesis, that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}\quad\text{for all }k=0,1,\dots,K-1.$ Hence, we get from (D)
$t_{k+1}\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq t_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+\rho_{k}t_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda}-\frac{1}{\gamma}\right)\|x_{k+1}-x_{k}\|^{2}$
$\quad+2\rho_{k}t_{k}(T_{k}+R_{k}).$ (40)
Now, for any $\rho_{k}\equiv\rho>0$, let us choose $\gamma$ such that $\gamma\leq(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda})^{-1}$, then the coefficient of $\|x_{k+1}-x_{k}\|^{2}$ in (D) is non-positive and may be discarded. Therefore, we have
$t_{k+1}\|\lambda_{k+1}-\lambda^{*}\|^{2}\leq t_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+2\rho t_{k}(T_{k}+R_{k}).$
Summing this inequality from $k=0$ to K1K-1 yields
$t_{K}\|\lambda_{K}-\lambda^{*}\|^{2}\leq t_{0}\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\sum_{k=0}^{K-1}t_{k}(T_{k}+R_{k}).$
Since $t_{k}\in[\frac{1}{2},1]$ for all $k\geq 0$, it follows that
$\frac{1}{2}\|\lambda_{K}-\lambda^{*}\|^{2}$ $\leq\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\sum_{k=0}^{K-1}(T_{k}+R_{k})$
$\leq\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\bigl(C_{T}+C_{R}(D_{\Lambda})\bigr),$
where the last inequality follows from the fact that $\sum_{k=0}^{K-1}T_{k}\leq C_{T}$ and $\sum_{k=0}^{K-1}R_{k}\leq C_{R}(D_{\Lambda}).$ Thus, we have
$\|\lambda_{K}-\lambda^{*}\|^{2}\leq 2\|\lambda_{0}-\lambda^{*}\|^{2}+4\rho\left(C_{T}+C_{R}(D_{\Lambda})\right)=2\|\lambda_{0}-\lambda^{*}\|^{2}+\frac{2\rho}{\gamma}\|x_{0}-x^{*}\|^{2}+8\rho L_{F,\theta}D_{\Theta}^{2}+4\rho C_{R}\left(D_{\Lambda}\right),$
where in the last equality we used the definition of $C_{T}$. Since both $\gamma=\mathcal{O}(D_{\Lambda})$ and $C_{R}\left(D_{\Lambda}\right)=\mathcal{O}(D_{\Lambda})$, selecting $D_{\Lambda}\triangleq B+\sqrt{A}$, where $A\triangleq 2\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\|x_{0}-x^{*}\|^{2}\bigl(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}\bigr)+8\rho L_{F,\theta}D_{\Theta}^{2}+4\rho\Bigg(\Bigl(L_{F,\theta}D_{X}+3L_{x\theta}^{\Phi}D_{X}+2J\rho D_{f}L_{f,\theta}+2\sqrt{J}L_{f,\theta}\|\lambda^{*}\|\Bigr)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+L_{F,\theta}\Bigl(D_{\Theta}^{2}+2\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}\Bigr)+\frac{L_{\lambda\theta}^{\Phi}}{2}\sum_{k=0}^{\infty}\frac{\|\theta_{k}-\theta^{*}\|^{2}}{\alpha_{k}}+\frac{J\rho}{2}L_{f,\theta}^{2}\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}\Bigg)$ and $B\triangleq 2\rho C_{2}\|x_{0}-x^{*}\|^{2}+8\rho\sqrt{J}\,L_{f,\theta}\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|$, implies that $\|\lambda_{K}-\lambda^{*}\|^{2}\leq D_{\Lambda}^{2}$ which completes the induction, hence, the result is proved. ∎
Appendix E Proof of Theorem 4.5
Proof.
Proof From (C), for every $k\geq 0$, $x\in X$, $\lambda\in\mathbb{R}^{J}_{+}$, we have
$(x_{k+1}-x)^{\top}F(x,\theta^{*})+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho}(x,\lambda_{k},\theta^{*})\leq\mathcal{T}_{k}(x)+\mathcal{R}_{k}(x,\lambda)+\frac{1}{2\rho}\bigl(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\bigr),$ (41)
where we grouped terms as follows
$\mathcal{T}_{k}(x)$ $\triangleq\frac{1}{2\gamma}\bigl(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\bigr)+\frac{L_{F,x}}{2}\big(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\big)$
$\quad+\big\langle r_{k+1},x_{k+1}-x\big\rangle-\big\langle r_{k},x_{k}-x\big\rangle,$
$\mathcal{R}_{k}(x,\lambda)$ $\triangleq\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\quad+\tfrac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}+\tfrac{L_{F,\theta}}{2}\big(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\big)+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|$
$\quad+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\big)+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
$\quad+\frac{J\rho}{2}L_{f,\theta}^{2}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$
Moving $\Phi_{\rho}(x,\lambda_{k},\theta^{*})$ to the right in (41), then summing over $k=0,\cdots,K-1$ and dividing by $K$, and using convexity of each $f_{j}(\cdot,\theta^{*})$ to pass the average inside, we get
$(\bar{x}_{K}-x)^{\top}F(x,\theta^{*})+f(\bar{x}_{K},\theta^{*})^{\top}\lambda$ $\leq\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})+\frac{1}{K}\sum_{k=0}^{K-1}\bigl(\mathcal{T}_{k}(x)+\mathcal{R}_{k}(x,\lambda)\bigr)$
$\quad+\frac{1}{2\rho}\bigl(\|\lambda_{0}-\lambda\|^{2}-\|\lambda_{K}-\lambda\|^{2}\bigr).$ (42)
Considering $\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)$, telescoping in the $x$-terms and the $r_{k}$-pair, we have
$\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)=\frac{1}{2\gamma}\left(\|x_{0}-x\|^{2}-\|x_{K}-x\|^{2}\right)-\frac{L_{F,x}}{2}\|x_{K}-x_{K-1}\|^{2}+\big\langle r_{K},x_{K}-x\big\rangle.$
Note that $\|r_{K}\|\leq L_{F,x}\|x_{K}-x_{K-1}\|+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|$, therefore $\big\langle r_{K},x_{K}-x\big\rangle\leq\frac{L_{F,x}}{2}\|x_{K}-x_{K-1}\|^{2}+\frac{L_{F,x}}{2}\|x_{K}-x\|^{2}+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|\|x_{K}-x\|$. Substituting this into the previous equality, we have
$\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)=\frac{1}{2\gamma}\|x_{0}-x\|^{2}+\left(\frac{L_{F,x}}{2}-\frac{1}{2\gamma}\right)\|x_{K}-x\|^{2}+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|\|x_{K}-x\|.$
Choosing $\gamma\in\left(0,L_{F,x}^{-1}\right)$, the coefficient of $\|x_{K}-x\|^{2}$ is non-positive and can be discarded. Since $x,x_{o},x_{K}\in X$ and $\theta_{K},\theta_{K-1}\in\Theta$, we have $\|x_{0}-x\|\leq D_{X},\quad\|x_{K}-x\|\leq D_{X}$, and $\|\theta_{K}-\theta_{K-1}\|\leq 2D_{\Theta}$. Therefore,
$\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)\leq\tfrac{1}{2\gamma}D_{X}^{2}+2L_{F,\theta}D_{X}D_{\Theta}\triangleq C^{\prime}_{T}\quad\forall K,\quad\forall x\in X.$ (43)
Next, we focus on $\sum_{k=0}^{K-1}\mathcal{R}_{k}(x,\lambda)$. By continuity on the compact sets $X,\Theta$ we have that $\sup_{x\in X,\theta\in\Theta}|f_{j}(x,\theta^{*})|\leq D_{f}$ for all $j=1,\cdots,J$. Now, let the constant-parameters $\rho_{k}\equiv\rho>0$ and $\gamma_{k}\equiv\gamma>0$. By Lemma 4.4, the dual sequence is bounded, and hence there exists $D_{\Lambda}>0$ such that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$ for all $k\geq 0.$ Therefore, $\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\,\|x_{k+1}-x_{k}\|^{2}\leq\frac{C_{2}D_{\Lambda}}{2}\|x_{k+1}-x_{k}\|^{2}$. Consequently, if the primal step-size is chosen so that $\gamma\in\Bigl(0,\,\bigl(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda}\bigr)^{-1}\Bigr),$ then the whole coefficient of $\|x_{k+1}-x_{k}\|^{2}$ in $\mathcal{R}_{k}(x,\lambda)$ is non-positive and may be discarded. The remaining terms in $\mathcal{R}_{k}(x,\lambda)$ are controlled by the boundedness of $X$, $\Theta$, and $\{\lambda_{k}\}_{k\geq 0}$ together with the learning-error summability assumption. In particular, since $\|\lambda_{k}\|\leq\|\lambda_{k}-\lambda^{*}\|+\|\lambda^{*}\|\leq D_{\Lambda}+\|\lambda^{*}\|$, the last term $2L_{f,\theta}(J\rho D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|$ is bounded by $2L_{f,\theta}\bigl(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\bigr)\|\theta_{k}-\theta^{*}\|$ which together with $\frac{J\rho}{2}L_{f,\theta}^{2}\|\theta_{k}-\theta^{*}\|^{2}$ are summable by the learning-error assumption. Therefore, we conclude that $\sum_{k=0}^{K-1}\mathcal{R}_{k}(x,\lambda)\leq C^{\prime}_{R}(x,\lambda)$ where
$C^{\prime}_{R}(x,\lambda)\triangleq$ $\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|$
$\quad+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}<+\infty.$ (44)
Using (43) and (E) in (E), we get
$(\bar{x}_{K}-x)^{\top}F(x,\theta^{*})+f(\bar{x}_{K},\theta^{*})^{\top}\lambda$
$\quad\leq\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})+\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,\lambda)}{K}+\frac{1}{2\rho K}\bigl(\|\lambda_{0}-\lambda\|^{2}-\|\lambda_{K}-\lambda\|^{2}\bigr).$ (45)
Let $x=x^{*}$ within (E). Since $x^{*}\in\mathcal{X}(\theta^{*})$, we have $\Phi_{\rho}(x^{*},\lambda_{k},\theta^{*})\leq 0$ for all $k$, hence
$(\bar{x}_{K}-x^{*})^{\top}F(x^{*},\theta^{*})+f(\bar{x}_{K},\theta^{*})^{\top}\lambda\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x^{*},\lambda)}{K}+\frac{1}{2\rho K}\bigl(\|\lambda_{0}-\lambda\|^{2}-\|\lambda_{K}-\lambda\|^{2}\bigr).$ (46)
By Lemma 2.2, evaluated at $(x^{*},\lambda^{*})$, we have
$(\bar{x}_{K}-x^{*})^{\top}F(x^{*},\theta^{*})+[f(\bar{x}_{K},\theta^{*})]_{+}^{\top}\lambda^{*}\geq 0.$ (47)
Subtracting (47) from (46) eliminates $(\bar{x}_{K}-x^{*})^{\top}F(x^{*},\theta^{*})$ and yields
$\Big(f(\bar{x}_{K},\theta^{*})^{\top}\lambda-[f(\bar{x}_{K},\theta^{*})]_{+}^{\top}\lambda^{*}\Big)\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x^{*},\lambda)}{K}+\frac{1}{2\rho K}\bigl(\|\lambda_{0}-\lambda\|^{2}-\|\lambda_{K}-\lambda\|^{2}\bigr),\quad\forall\lambda\in\mathbb{R}_{+}^{J}.$ (48)
Let $\lambda=\tilde{\lambda}$ within (48) such that $\tilde{\lambda}_{j}\triangleq\begin{cases}1+\lambda_{j}^{*},&\text{if }f_{j}(\bar{x}_{K},\theta^{*})>0\\ 0,&\text{otherwise}\end{cases}$, hence, $f(\bar{x}_{K},\theta^{*})^{\top}\tilde{\lambda}-[f(\bar{x}_{K},\theta^{*})]_{+}^{\top}\lambda^{*}=\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}$ which implies that
$\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x^{*},\tilde{\lambda})}{K}+\frac{J+\|\lambda_{0}-\lambda^{*}\|^{2}}{\rho K}=\frac{C_{\mathrm{feas}}}{K},$ (49)
where $C_{\mathrm{feas}}=C^{\prime}_{T}+\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J+\|\lambda_{0}-\lambda^{*}\|^{2}}{\rho}$.
Moreover, if we define $\epsilon\triangleq\frac{C_{\mathrm{feas}}}{K}$ and the enlarged feasible set $\mathcal{X}_{\epsilon}(\theta^{*})=\{x\in X\mid\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon\}$. Then, $\sum_{j=1}^{J}[f_{j}(x,\theta^{*})]_{+}\leq\epsilon$, and using the definition of $\Phi_{\rho}$ together with $\|\lambda_{k}\|\leq D_{\Lambda}$ yields
$\Phi_{\rho}(x,\lambda_{k},\theta^{*})$ $=\sum_{j=1}^{J}\frac{\bigl([\lambda_{k}^{(j)}+\rho f_{j}(x,\theta^{*})]_{+}\bigr)^{2}-\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho}$
$\leq\sum_{j=1}^{J}\Bigl(\lambda_{k}^{(j)}[f_{j}(x,\theta^{*})]_{+}+\frac{\rho}{2}[f_{j}(x,\theta^{*})]_{+}^{2}\Bigr)$
$\leq D_{\Lambda}\,\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}+\frac{\rho}{2}\bigl(\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\bigr)^{2}$
$\leq D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}.$
Note that this bound is uniform in $k$. Therefore, we have $\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})\leq D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}.$ Now, let $\lambda=\mathbf{0}$ in (E) to conclude that
$(\bar{x}_{K}-x)^{\top}F(x,\theta^{*})$ $\leq\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})+\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}$
$\leq D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}+\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}$
$\leq D_{\Lambda}\frac{C_{\mathrm{feas}}}{K}+\frac{\rho}{2}\left(\frac{C_{\mathrm{feas}}}{K}\right)^{2}+\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}.$ (50)
Taking the supremum over $x\in\mathcal{X}_{\epsilon_{K}}(\theta^{*})$ gives
$\widetilde{\mathrm{Gap}}(\bar{x}_{K},\theta^{*})\triangleq\sup_{x\in\mathcal{X}_{\epsilon_{K}}(\theta^{*})}F(x,\theta^{*})^{\top}(\bar{x}_{K}-x)\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}+D_{\Lambda}\frac{C_{\mathrm{feas}}}{K}+\frac{\rho}{2}\left(\frac{C_{\mathrm{feas}}}{K}\right)^{2},$ (51)
where $C^{\prime}_{T}$, and $C_{\mathrm{feas}}$ are defined as above and $C^{\prime}_{R}(x,0)=\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}.$